OpenStudy (anonymous):

a/b = c/d.. is it proportional to: a. a/c=b/d? (yes or no) b. d/b=c/a? (yes or no) c. c/b=a/d? (yes or no) d. a/d=c/b? (yes or no)

ganeshie8 (ganeshie8):

cross multiply each of the proportions, and see which are same

ganeshie8 (ganeshie8):

a/b = c/d => ad = bc

ganeshie8 (ganeshie8):

next do the same for given options, and see which ones are the same

ganeshie8 (ganeshie8):

a. a/c=b/d? (yes or no) cross multiplying :- ad = bc so YES

ganeshie8 (ganeshie8):

there is one more option, see if u can find it

OpenStudy (anonymous):

So, after cross multiplying, the both sides must be equal ?:) @ganeshie8

ganeshie8 (ganeshie8):

nope, after cross multiplicaiton u should get "ad = bc" which is the GIVEN.

ganeshie8 (ganeshie8):

try option b, and tell me wat u get after cross multiplying

OpenStudy (anonymous):

Oh, alright, I had answers but I wasn't so sure. Thanks :)

OpenStudy (anonymous):

what if its.. \[\frac{ a+b }{ b } = \frac{ c+d }{ d }\] ? It will be.. (a+b)(d) = (c+d)(b) and what's the next step? @ganeshie8

ganeshie8 (ganeshie8):

wat you trying to do here ?

ganeshie8 (ganeshie8):

(a+b)(d) = (c+d)(b) ad + bd = bc + bd ad = bc

ganeshie8 (ganeshie8):

so yes, this also is same as given proportion

OpenStudy (anonymous):

oh thanks ;)

ganeshie8 (ganeshie8):

you may also look at it this way :- \(\large \frac{ a+b }{ b } = \frac{ c+d }{ d }\) \(\large \frac{ a }{ b } + \frac{ b }{ b } = \frac{ c }{ d } + \frac{ d }{ d } \) \(\large \frac{ a }{ b } + 1 = \frac{ c }{ d } + 1\) \(\large \frac{ a }{ b } = \frac{ c }{ d } \) which is same as given proportion !

OpenStudy (anonymous):

oh yeah, so informative, thanks again :DD

ganeshie8 (ganeshie8):

np :)